OpenStudy (anonymous):

need help with solving a equation with substitution I want it explained so i can understand, this is the equation; x + 3y = 56

OpenStudy (anonymous):

What are you substituting?

OpenStudy (anonymous):

there is 56 houses with two floorplans, y is 3 times more than x floorplan

OpenStudy (anonymous):

y=3x

OpenStudy (anonymous):

Is there another question to help you out in solving this? Substitution is when you take one equation and plug it into another. You can do this by solving one of the usually 2 equations for either x or y. Once you have one of those 2 variables isolated you plug it back into the other equation to solve for the variable. Once you find that variable you plug it back into the equation you substitued with to get the other variable... for example if you have: 2x + y = 3 x + y = 1 Take the x + y = 1 and solve for y. You will get y = 1-x. Then with that in mind you plug that y back into the 2x +y = 3. So you will get 2x + 1 - x = 3. Then you simplify and solve for x. ONce you find x you plug it into y = 1-xto find y.

OpenStudy (anonymous):

x+3(3x)=56 4x=56 x=14

OpenStudy (anonymous):

Sorry I was called out for a moment, abigail2010, if you are still there, I only have one question, that is why I do not understand using substitution. My problem is, there is 56 houses for sale with two floor plans. I am to use x to represent floor plan one, and y for floor plan two. Then there is 3 times as many homes available with floor plan two, and I am to figure out how many of each floor plans are avaailable using the substitution